OpenStudy (anonymous):

if Z= eˆx (x cos y - y sin y) show that ∂²z/∂x² + ∂²z/∂y² = 0

OpenStudy (anonymous):

you have to compute right? \[Z_x=e^x\cos(y)+e^x(x\cos(y)-y\sin(y))\] \[Z_{xx}=\cos(y)e^x+e^x\cos(y)+e^x(x\cos(y)-y\sin(y))\] unless i made a mistake. clean it up with some algebra, then repeat for \(Z_{yy}\) and it should work out

OpenStudy (anonymous):

Because of all of the power the colonists had due to their usage of gun powder, military knowledge and even their god given right to commit genocide and invade foreign land in order to steal their resources. The cause of the problem was the greed the Europeans had for precious materials and resources to keep their empires at bay. I hold the colonists responsible for leading his people into committing atrocious genoside against the natives. One of the many problems that have resulted as a consequence of ignoring the genocide every thanksgiving is the unresolved anger that has built up throught the years by the victims of such atrocities. We can reach a more detailed account of these events in the book A People’s History by Howard Zinn. To correct the problem, the colonists should have used logic to have the Natives as friends to help them find their precious materials.